A galvanometer with a resistance of 15.0Ω gives a full-scale reading for current of 3.0 A. What shunt resistance is needed to convert this into an ammeter with a 30.Ω A full-scale reading?

1
Expert's answer
2022-02-20T15:46:35-0500

We know that

"S=\frac{i_g G}{I-i_g}"

"i_g=\frac{SI}{S+G}"

"i_g=\frac{SI}{S+G}"

Put value

"i_g=\frac{15\times3}{30+15}=1A"